An 18 foot ladder is leaning against a house. The bottom of the ladder is sitting on flat ground and is 7 ft away from bottom of house. What is the height where the top of the ladder is touching the house?
Answer by rfer(16322)

18^2-7^2=
324-49=275
sq rt 275=16.58 ft